# AgentNoiseBench: Benchmarking Robustness of Tool-Using LLM Agents Under Noisy Condition

### v1

@@ 2.1 Design Principles
Trajectory Consistency Enforcement. Outcome-based metrics are insufficient as they fail to distinguish between robust reasoning and stochastic luck. Therefore, a rigorous evaluation must demand procedural integrity: the agent’s intermediate reasoning steps should remain logically sound and strictly adhere to task requirements, impervious to noise-induced deviations. This ensures that success reflects genuine stability in the decision-making process rather than accidental correctness.
@@ 2.2 Benchmark Construction
To ensure a fair and standardized comparison across different agents, our benchmark construction adheres to a unified testbed principle. We employ a data-driven analysis of user–agent interaction logs to derive realistic noise patterns. Furthermore, to guarantee that adversarial challenges remain consistent across all evaluated models, we adopt a proxy-guided generation strategy, ensuring that the difficulty level is controlled by a fixed, standardized noise generation protocol rather than adapting to specific victim models.
@@ 2.2.1 Empirical Taxonomy of Noise
We derive a noise taxonomy through systematic analysis of large-scale user–agent interaction logs. Specifically, we combine manual annotation with frequency-based statistics to identify noise patterns that are both prevalent in practice and associated with high interaction costs.
@@
User noise captures imperfections and variability in human-provided instructions during agent interactions. Specifically, we consider the following categories of user-side perturbations:
Ambiguous Instructions: Instructions that are underspecified or lack explicit constraints.
Inconsistent Instructions: Contradictory or conflicting directives issued across multi-turn interactions.
Redundant Information: The inclusion of excessive irrelevant or low-utility contextual information.
Topic Drift: Shifts in the user’s primary intent or conversational focus across dialogue turns.
Boundary Probing: Queries that deliberately test the limits of the agent’s knowledge or reasoning capabilities.
@@
Tool noise captures the stochasticity and unreliability of external tools, APIs, and environments during task execution. Specifically, we consider the following categories of tool-side perturbations:
@@
Incomplete Responses: Tools return partial or malformed outputs, such as truncated JSON objects or missing fields required for downstream steps.
Erroneous Outputs: Tools produce responses containing incorrect or internally inconsistent factual information.
@@ 2.2.2 Constrained Adversarial Noise Injection
To generate noise that is strictly valid yet sufficiently challenging, while ensuring fairness across diverse models, we introduce a constrained adversarial injection mechanism. Specifically, we employ a state-of-the-art model (e.g., o3) as a fixed reference agent to derive a universal noise generation system prompt.
We instantiate a strong language model as the noise generator , with its parameters frozen. Optimization is performed exclusively over the parameterized adversarial system prompt . Given an original task input , the generator produces a perturbed task instance conditioned on . Our objective is to identify an optimal prompt strategy that maximally degrades the performance of the reference agent , subject to the strict preservation of task solvability:
Here, quantifies the performance degradation of the reference agent under injected noise, while enforces a solvability constraint to exclude ill-posed task instances. We iteratively update to steer the model toward producing adversarial perturbations that induce maximal failure without violating task validity.
By anchoring the optimization process to a state-of-the-art proxy, we leverage the transferability of adversarial examples: noise capable of disrupting the most robust agent serves as a rigorous, high-bar challenge for all other models. Once identified, is frozen and applied uniformly to generate noise for all evaluated agents. This ensures that the benchmark difficulty remains static and unbiased, preventing adaptive overfitting to specific victim models, thereby systematically constructing the hardest valid evaluation instances.
@@ 2.3 Trajectory-Aware Evaluation Protocol
We retain the standard evaluation metrics from VitaBench (He et al., 2025) and -Bench (Barres et al., 2025) to assess final outcome correctness. To rigorously capture procedural integrity, we additionally introduce a Trajectory-Aware Evaluation Protocol. This mechanism scrutinizes the entire interaction history, penalizing agents that arrive at correct answers via corrupted or unsafe reasoning paths.
Trajectory Deviation Detection. Let denote the agent’s reasoning trajectory. Given a task specification and a noise injection plan , we first define a step-wise validity indicator . For each state , a model-based evaluator verifies whether the agent’s behavior remains consistent with the task requirements:
A trajectory is deemed valid only if all intermediate steps maintain consistency. Consequently, the trajectory-level validity is defined as the conjunction of step-wise indicators:
Stability-Gated Success Metric. To strictly enforce robustness, we propose a stability-gated accuracy (SGA) metric. This metric regards a task as successfully solved if and only if the agent produces the correct outcome and maintains a flawless reasoning trajectory:
where denotes the binary success of the final outcome. Under this rigorous standard, any intermediate deviation () collapses the entire metric to zero, effectively filtering out false positive success cases driven by stochastic luck.
@@ 3.1 Experimental Setup
Datasets. To systematically evaluate agent robustness in complex environments, we delineate core capabilities into two dimensions: (1) static knowledge retrieval and (2) dynamic multi-turn interaction. Specifically, for static retrieval tasks, we employ two representative benchmarks: 2WikiMultiHopQA (Ho et al., 2020) and HotpotQA (Yang et al., 2018). Collectively, the averaged performance across these two benchmarks is reported as our ’Search’ metric, representing static knowledge retrieval. Given that these tasks inherently lack user interaction and solely involve tool calls, we primarily introduce tool noise. Conversely, for dynamic interaction tasks, we utilize VitaBench and -Bench, two benchmarks that support the simultaneous injection of both user noise and tool noise. In each dimension, noisy evaluation instances are constructed by perturbing the original benchmark tasks as base datasets. Across all test domains, noise is introduced by perturbing key interaction interfaces, while strictly preserving the invariance of the original task objectives. Finally, to strike a balance between the experimental necessity of covering all predefined noise types and the practical constraints of evaluation scalability and cost, we adopt a stratified sampling strategy: 25% of samples are drawn from each scenario within VitaBench and -Bench, and 500 instances are sampled from the multi-hop QA datasets. This design ensures precise comparative analysis of model performance in both noisy and noise-free environments under a unified task distribution.
Metrics. We report mean performance over independent trials per task. Our evaluation captures both effectiveness and efficiency: Avg@4: The average task success rate across trials, serving as the primary measure of agent robustness (effectiveness). Avg_Tokens@4: The average token consumption across trials, quantifying computational cos Avg_Steps@4: The average number of interaction steps required to complete the task across trials, quantifying computational cost. Robustness: The relative rate of change of a metric above from noise-free to noisy environments, quantifying agent robustness under noisy conditions.
@@ 3.2 Noise Robustness Remains an Open Challenge for Current Agents
Observation 1: Agents universally exhibit significant vulnerability to noise, with varying degrees of sensitivity. Experimental result reveals that nearly all models suffer severe performance degradation upon the injection of noise, with an average accuracy drop of up to 20.8%. This performance loss is prevalent across all evaluation domains; for instance, the average declines in the Vitbench, -bench, and Search domains reach 24.1%, 21.2%, and 16.8%, respectively. Furthermore, we observe significant disparities in the magnitude of performance degradation across different models when subjected to identical noise interference. These results not only reveal the significant performance gap between an ideal noise-free environment and a real noisy scenario, but also highlight the intrinsic differences in models’ sensitivity to noise, posing a serious challenge to the widespread deployment of current agents.
@@ 3.3 Misalignment Exists between Reasoning Ability and Robustness
Observation 2: Reasoning ability and robustness are not closely related. Our experimental results indicate that reasoning ability and noise robustness are not closely related. Specifically, as shown in Figure 3, with the exception of specific scenarios (e.g., Delivery), the robustness scores of thinking models are relatively lower than those of non-thinking models across most other scenarios. This performance disadvantage prevalent across multiple domains clearly indicates that strong reasoning ability do not inherently confer robustness against noise. To deeply investigate the root causes of the low robustness observed in powerful reasoning models under noisy conditions, we analyzed the models’ reasoning trajectorys. Our study uncovered an interesting phenomenon: reasoning models often fail to effectively filter out noisy information. Instead, they tend to misinterpret this noise as meaningful, critical signals, and on this basis, attempt to construct complex logical chains to rationalize the noise. This mechanism leads to the generation of spurious reasoning chains—where the model produces solutions that appear logically rigorous but are in fact incorrect. The fundamental reason for this behavior is closely related to the training of reasoning models; their training data is typically strictly screened and pre-processed to eliminate noise. Consequently, when noise appears during the testing phase, the powerful reasoning capability of these models more readily fabricates information to make the noise seem rational.
@@ 3.4 How Noise Disrupts Reasoning via Entropy Injection
Observation 3: Noise alter entropy via distinct mechanisms. The results indicate that different noise sources alter entropy changes in the reasoning process through fundamentally different mechanisms. In noise-free baselines, we observe entropy decay, where the agent progressively resolves uncertainty, converging towards a deterministic solution. In contrast, user noise induces a high intercept entropy profile, characterized by extreme uncertainty at the initial step. This reflects the cognitive load required to disambiguate user intent before any planning can commence. Crucially, as shown in Figure 4, tool noise results in a relatively high level of entropy starting from the middle stage and persisting through the late stage of the agent trajectory. Here, the output of hallucination-inducing tools introduces uncertain information into the context, making it difficult for the agent to establish a reliable chain of reasoning.
3.5 Significant Performance Disparities Induced by Noise from Different Sources and Granularities
To comprehensively characterize the distinct impact of noise on agent performance, we conduct a two-level analysis from both the macroscopic scenario and the microscopic noise granularity, as shown in Figure 5(b) and Figure 5(a), respectively. Based on the experimental results, we summarize the following observations:
Observation 4: Fine-grained noise types exhibit markedly different destructive effects on agent performance. Our results demonstrate that the impact of noise is highly non-uniform and strongly depends on the specific noise type.
On the tool side, execution failures are the most destructive, reducing the average performance score to 0.31. This severe degradation arises because such noise induces structural execution breakdowns, which sever the agent’s access to environmental feedback. As a result, the reasoning process can no longer progress due to the absence of factual grounding. In contrast, noise types such as incomplete information or redundancy exhibit relatively milder effects. This suggests that agents retain a limited ability to filter and complete partial information, enabling them to extract useful signals from noisy tool outputs and sustain basic decision-making chains.
On the user side, instruction conflicts are the most harmful. When confronted with logically contradictory instructions, agents struggle to infer a consistent and correct intent, leading to pronounced performance degradation. By contrast, instruction ambiguity causes only minor performance drops. This indicates that although ambiguous intent increases interpretation difficulty, agents can often rely on the strong prior knowledge embedded in large language models to perform probabilistic semantic repair, thereby partially mitigating the adverse impact of noise.
Observation 5: Agents exhibit substantially higher sensitivity to tool-side noise than to user-side noise. Empirical results reveal a pronounced robustness gap between the two noise sources: across all agent families, vulnerability to tool-side noise is significantly higher than to user-side noise. Quantitatively, tool noise typically leads to severe performance drops of 20%–30%, whereas user noise generally results in reductions of around 10
To explain this asymmetric sensitivity, we further analyze the agent–environment interaction process and identify a fundamental distinction between user noise and environmental (tool) noise during reasoning.
@@ 3.5 Significant Performance Disparities Induced by Noise from Different Sources and Granularities
User noise is semantically repairable. User noise primarily manifests as perturbations at the linguistic level. During pretraining, large language models acquire extensive linguistic and world knowledge, endowing them with strong error-correction and associative capabilities. Consequently, even when user instructions are noisy, agents can often internally repair these perturbations at the semantic level, allowing task execution to proceed.
@@
Tool noise causes a breakdown of factual grounding. In contrast, tool-side noise directly corrupts the agent’s only channel for perceiving the external world. During task execution, tool outputs serve as the sole objective evidence for state estimation and action selection. Once this factual grounding is disrupted, the loss is structural rather than semantic: agents cannot infer the true environment state through internal reasoning alone. As a result, the reasoning chain loses its support and collapses, ultimately leading to task failure.
3.6 When Does Noise Injection Do the Most Impact
@@
To investigate how the timing of noise injection affects agent performance, we consider three injection stages: early, middle, and late. We conduct experiments on four representative model families, including Claude, Gemini, DeepSeek, and LongCat. The results are shown in Figure 6, with additional results for more models reported in Appendix A.2. Based on these results, we derive the following observation:
Observation 6: Agents are most sensitive to noise injected at the middle stage. As illustrated in Figure 6, across all four models, injecting noise during the middle stage leads to a significantly larger performance degradation than injecting noise at either the early or late stage. This indicates that agents exhibit relatively higher robustness to noise introduced at the early or the late of the trajectory, while being more sensitive to noise during the middle stage. Importantly, the consistency of this trend across models with different architectures and reasoning strategies suggests that the heightened sensitivity during the intermediate stage of the process may be an inherent property, rather than a model-specific phenomenon.

### v2

@@ 2.1 Design Principles
Realistic Evaluation Requirements. Outcome-based metrics are insufficient as they fail to distinguish between robust reasoning and stochastic luck. Therefore, a rigorous evaluation must demand process-level correctness: the agent’s intermediate reasoning steps should remain logically sound and strictly adhere to task requirements, impervious to noise-induced deviations. This ensures that the evaluation metrics faithfully reflect the agent’s robustness to noise.
@@ 2.2 Benchmark Construction
To enable fair comparisons across agents, AgentNoiseBench adopts a standardized benchmark construction pipeline. Concretely, we first derive realistic noise types from large-scale user–agent interaction logs. We then instantiate a fixed adversarial noise generator to produce challenging yet valid perturbations, and apply it uniformly to all evaluated agents. This design keeps the noise distribution and difficulty level consistent across models, thereby enabling controlled and fair evaluations.
@@ 2.2.1 Empirical Taxonomy of Noise
We categorize noise types through a systematic analysis of large-scale user–agent interaction logs. Concretely, we combine manual annotation with frequency-based statistical analysis to identify noise types that are both prevalent in real-world interactions and associated with high interaction costs. We group the resulting noise into two broad families: user noise and tool noise.
@@
Models the imperfections and variability in human-provided instructions during agent interactions. Concretely, we consider the following categories of user-side noise:
Ambiguous Instructions: User instructions that are underspecified or lack explicit constraints.
Inconsistent Instructions: User instructions are contradictory or conflicting across multi-turn interactions.
Redundant Information: User instruction include irrelevant contextual information.
Topic Drift: User instructions that shift the primary task intent or conversational focus across dialogue turns.
Boundary Probing: User instructions that deliberately test the limits of the agent’s knowledge.
@@
Models the stochasticity and unreliability of external tools, APIs, and environments that the agent interacts with. Concretely, we consider the following categories of tool-side noise:
@@
Incomplete Responses: Tools return outputs with missing or partially truncated data.
Erroneous Outputs: Tool responses containing incorrect or inconsistent factual information.
@@ 2.2.2 Constrained Adversarial Noise Injection
To generate noise that is strictly valid yet sufficiently challenging, while ensuring fairness across diverse models, we introduce a constrained adversarial injection mechanism. Specifically, we employ a state-of-the-art model (e.g., o3) as a fixed reference agent to construct a universal noise generation system prompt.
We instantiate a strong language model as the noise generator , with its parameters frozen. Optimization is performed exclusively over the adversarial system prompt . Given an agent input at a particular interaction step, which is either a user instruction or a tool return, the generator produces a perturbed input . Our objective is to identify an optimal prompt strategy that maximally degrades the performance of the reference agent , subject to strict preservation of the underlying task solvability:
Here, quantifies the performance degradation of the reference agent under injected noise, while enforces a constraint ensuring that the perturbed agent input remains solvable. We iteratively update to maximize under the solvability constraint, thereby producing valid yet adversarial noise.
By optimizing the noise prompt against a state-of-the-art reference agent, we leverage the transferability of noise across agents: noise that reliably disrupts a highly capable reference agent tends to generalize and remain challenging for other agents. After obtaining , we freeze it and apply it uniformly when generating noise for all evaluated agents. This ensures that all agents are evaluated under identical noise conditions, enabling a fair and consistent comparison.
@@ 2.3 Trajectory-Aware Evaluation Protocol
We retain the standard evaluation metrics from VitaBench (He et al., 2025) and -Bench (Barres et al., 2025) to assess final outcome correctness. To more comprehensively assess agent behavior, we additionally introduce a Trajectory-Aware Evaluation Protocol. This protocol examines the full interaction trajectory and accounts for cases where a correct final answer is obtained via deviated reasoning or interaction paths.
Trajectory Deviation Detection. Let denote the agent’s reasoning trajectory. Given task requirements , we define a step-wise validity indicator . For each state , assesses whether the agent’s behavior at that step exhibits a noise-induced deviation:
A trajectory is deemed valid only if all steps remain consistent. Consequently, the trajectory-level validity is defined as the conjunction of step-wise indicators:
Stability-Gated Success Criterion. We apply trajectory-level validity as a gating condition for task execution success. Under this criterion, a task is considered successfully solved only when the agent produces the correct final outcome and the corresponding reasoning trajectory remains valid throughout the interaction:
where indicates whether the final outcome is correct for task . This criterion filters out cases where a correct final outcome is achieved despite deviations in the interaction trajectory.
@@ 3.1 Experimental Setup
Datasets. To systematically evaluate agent robustness in complex environments, we delineate core capabilities into two dimensions: (1) static knowledge retrieval and (2) dynamic multi-turn interaction. Specifically, for static retrieval tasks, we employ two representative benchmarks: 2WikiMultiHopQA (Ho et al., 2020) and HotpotQA (Yang et al., 2018). Collectively, the averaged performance across these two benchmarks is reported as ’Search’ , representing static knowledge retrieval. Given that these tasks inherently lack user interaction, tool noise is injected. Conversely, for dynamic interaction tasks, we utilize VitaBench and -Bench, two benchmarks that support the simultaneous injection of both user noise and tool noise. Across all test domains, noise is introduced by perturbing key interaction interfaces, while strictly preserving the invariance of the original task objectives. Finally, to strike a balance between the experimental necessity of covering all predefined noise types and the practical constraints of evaluation scalability and cost, we adopt a stratified sampling strategy: 25% of samples are drawn from each scenario within VitaBench and -Bench, and 500 instances are sampled from the multi-hop QA datasets. This design ensures precise comparative analysis of model performance in both noisy and noise-free environments under a unified task distribution.
Metrics. We report mean performance over independent trials per task. Our evaluation captures both effectiveness and efficiency: Avg@4: The average task success rate across trials, serving as the primary measure of agent robustness. Avg_Tokens@4: The average token consumption across trials, quantifying computational cos Avg_Steps@4: The average number of interaction steps required to complete the task across trials, quantifying computational cost. Robustness: The relative rate of change of a metric above from noise-free to noisy environments, quantifying agent robustness under noisy conditions.
@@ 3.2 Noise Robustness Remains an Open Challenge for Current Agents
Observation 1: Agents universally exhibit significant vulnerability to noise, with varying degrees of sensitivity. Our results indicate that nearly all models suffer substantial performance degradation upon the injection of noise, with an average accuracy drop of 20.8%. This performance loss is prevalent across all evaluation domains; for instance, the average declines in the Vitbench, -bench, and Search domains reach 24.1%, 21.2%, and 16.8%, respectively. Furthermore, we observe significant disparities in the magnitude of performance degradation across different models when subjected to identical noise interference. These results not only reveal the significant performance gap between an ideal noise-free environment and a real noisy scenario, but also highlight the intrinsic differences in models’ sensitivity to noise, raising concerns for the reliable deployment of current agents.
@@ 3.3 Misalignment Exists between Reasoning Ability and Robustness
Observation 2: Reasoning ability and robustness are not closely related. Our experimental results indicate that reasoning ability and noise robustness are not closely related. Specifically, as shown in Figure 3, with the exception of specific scenarios (e.g., Delivery), the robustness scores of thinking models are relatively lower than those of non-thinking models across most other scenarios. This performance disadvantage prevalent across multiple domains clearly indicates that strong reasoning ability does not inherently confer robustness against noise. To deeply investigate the root causes of the low robustness observed in powerful reasoning models under noisy conditions, we analyzed the models’ reasoning trajectories. Our study uncovered an interesting phenomenon: reasoning models often fail to effectively filter out noisy information. Instead, they tend to misinterpret this noise as meaningful, critical signals, and on this basis, attempt to construct complex logical chains to rationalize the noise. This mechanism leads to the generation of spurious reasoning chains—where the model produces solutions that appear logically rigorous but are in fact incorrect. The fundamental reason for this behavior is closely related to the training of reasoning models; their training data is typically strictly screened and pre-processed to eliminate noise. Consequently, when noise appears during the testing phase, the powerful reasoning capability of these models more readily fabricates information to make the noise seem rational.
@@ 3.4 How Noise Disrupts Reasoning via Entropy Injection
Observation 3: Noise perturbs reasoning entropy through distinct mechanisms. The results indicate that different noise sources alter entropy changes in the reasoning process through fundamentally different mechanisms. In noise-free baselines, we observe entropy decay, where the agent progressively resolves uncertainty, converging towards a deterministic solution. In contrast, user noise induces a high intercept entropy profile, characterized by extreme uncertainty at the initial step. This reflects the cognitive load required to disambiguate user intent before any planning can commence. Crucially, as shown in Figure 4, tool noise results in a relatively high level of entropy starting from the middle stage and persisting through the late stage of the agent trajectory. Here, the output of hallucination-inducing tools introduces uncertain information into the context, making it difficult for the agent to establish a reliable chain of reasoning.
3.5 Impact of Noise Sources and Granularity on Agent Performance
To further investigate the distinct effects of noise on agents, as shown in Figure 5(b) and Figure 5(a), we analyze noise from both a macro perspective of noise sources and a micro perspective of fine-grained noise types. Based on the experimental results, we summarize the following observations.
Observation 4: Agents exhibit higher sensitivity to tool-side noise. The results reveal a substantial robustness gap when agents are exposed to different noise sources: across model families and scenarios, agents are significantly more sensitive to tool noise than to user noise. Quantitative analysis shows that tool noise typically results in severe performance degradation of 20To explain this asymmetric sensitivity, we analyze the agent interaction logic and attribute it to differences in the nature of user noise and tool noise.
@@ 3.5 Impact of Noise Sources and Granularity on Agent Performance
User noise exhibits semantic repairability. User noise is essentially a perturbation at the level of language expression. During pretraining, large language models acquire extensive prior knowledge, which endows them with strong error correction and associative capabilities. As a result, even when input instructions are noisy, agents can often leverage internal commonsense knowledge to repair such noise at the semantic level, thereby maintaining normal task execution.
@@
In contrast, tool noise leads to a complete loss of reasoning evidence. Tool-side noise directly corrupts the agent’s channel for perceiving the external environment. During task execution, tool outputs serve as the only objective basis for assessing the current state and deciding subsequent actions. The absence or corruption of such objective facts constitutes a structural failure: agents cannot infer the environment state from internal reasoning, causing the reasoning chain to lose its foundation and ultimately leading to task failure.
Observation 5: Fine-grained noise types exhibit significantly different destructive effects on agents. The experimental results indicate that the impact of fine-grained noise on agents’ performance is highly uneven. On the user side, instruction contradiction is the most destructive noise type. Agents struggle to infer a consistent and correct intent from logically incompatible instructions, resulting in a pronounced performance drop. In comparison, instruction ambiguity leads to only limited performance degradation. This suggests that although ambiguous intent increases interpretation difficulty, agents can often rely on probabilistic intent inference and partial semantic repair, supported by the rich prior knowledge of large language models, thereby partially mitigating the negative impact of noise. On the tool side, execution failures are the most destructive, reducing the average performance score to 0.31. A reasonable explanation is that such noise triggers structural interruptions in execution, cutting off the agent’s access to environmental feedback and preventing the reasoning process from progressing due to the lack of external factual grounding. By contrast, noise such as incomplete information or redundancy have relatively milder effects, suggesting that agents retain some ability to filter and complete information, extracting useful signals from noisy feedback to sustain correct reasoning processes.
3.6 When Does Noise Injection Have the Greatest Impact
@@
To investigate the impact of noise injection timing on agent performance, we divide the task execution trajectory into three noise injection stages: early stage, middle stage, and late stage. We conduct experiments on four representative model families, Claude, Gemini, DeepSeek, and LongCat. The experimental results are shown in Figure 6, with additional results for more models provided in Appendix A.2. Based on these results, we summarize the following observation:
Observation 6: Agents are most sensitive to noise injected at the middle stage. Specifically, for the majority of evaluated models, noise injected at the middle stage leads to significantly greater performance degradation than noise injected at either the early or late stage, indicating that agents are more sensitive to perturbations occurring in the middle of the execution trajectory. One possible explanation is that, at the early stage, the model maintains higher uncertainty about the environment and thus reasons more conservatively, while at the late stage the state space has largely contracted and action choices are more constrained. In contrast, during the middle stage, the model tends to make more deterministic decisions based on accumulated contextual information, causing injected noise to be more easily absorbed as reliable signals. Notably, this phenomenon is highly consistent across models with different architectures and reasoning strategies, suggesting that heightened sensitivity to middle-stage noise is likely an inherent property of the agent’s execution process rather than a model-specific phenomenon.